# On Signed Network Coordination Games

### v1

@@ I Introduction
Network games comprising both coordinating and anti-coordinating interactions have been recently proposed to model the presence of anti-conformist behaviors in a social community, accounting for some form of heterogeneity [22, 23, 24, 25] . More generally, games exhibiting both strategic complements and substitutes have been recently proposed in the economic literature [26] to model heterogeneous interactions, e.g., markets with coexistence of both Cournot and Bertrand type firms. Such mixed games may fail to possess Nash equilibria. A fundamental example is the matching pennies game, which is a two-player game with one coordinating and one anti-coordinating player. In contrast with the matching pennies game, one can imagine that in a scenario where most of the players are coordinating and form a ’well’ connected subset, the presence of few anti-coordinating players should not prevent the coordinating players to reach a consensus and possibly the whole system to reach a Nash equilibrium. This is one main motivation to our work.
@@ II-A Signed Network Coordination Games
Figure 1 illustrates a network with weight values reported next to the corresponding links. Notice that, for =$\mathcal{R}=\{1,\dots,8\}$ and ~=$\tilde{\mathcal{R}}=\{1,\dots,6\}$, the subnetworks $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\tilde{\mathcal{R}}}$ are both unsigned, whereas for $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}=\{9,\dots,13\}$ and $\tilde{\mathcal{S}}=\mathcal{V}\setminus\tilde{\mathcal{R}}=\{7,\dots,13\}$ the subnetworks $\mathcal{G}_{\mathcal{S}}$ and $\mathcal{G}_{\tilde{\mathcal{S}}}$ are both undirected.
@@
We now provide an interpretation of the SNC games introduced above. The utility function (1) is the sum of two terms. The first addend in the right-hand side of (1) is a standalone term $h_{i}x_{i}$ (i.e., a term that depends only on the action of player $i$) that models the bias of player $i$ towards one of the two possible actions depending on the sign of the preference term $h_{i}$: if $h_{i}>0$ (respectively, $h_{i}<0$), then player $i$ has a bias towards action $+1$ ($-1$), in that, in the absence of the network (i.e., if $W=0$) this action would provide a utility that is $2h_{i}$ larger that its alternative $-1$ ($+1$); if $h_{i}=0$, then no bias is present in that both actions $+1$ and $-1$ would reward player $i$ with the same utility in the absence of the network. The second addend in the right-hand side of (1) is the aggregate of pairwise interaction terms $x_{i}W_{ij}x_{j}$, each of which can be interpreted as the utility of player $i$ in a two-player game with player $j$. Notice that, while the sum index $j$ in the right-hand side of (1) formally runs over all players $j$ in $\mathcal{V}$, the sum is effectively only over players $j$ such that $W_{ij}\neq 0$, i.e., out-neighbors of player $i$ in the network $\mathcal{G}$. Observe that a SNC game with binary actions is an instance of a polymatrix or pairwise-separable game. [41, 42].
@@ II-B Best Responses and Nash Equilibria
The best response (BR) correspondence for a player $i$ in $\mathcal{V}$ is the set of its optimal actions given the strategy profile or the other players. Formally,
@@
The notions of BR-reachability, BR-invariance, and BR-stability are particularly relevant for the study of so-called asynchronous best response dynamics [43]. These are a class of asynchronous dynamics modeled as discrete-time Markov chains with finite state space coinciding with the strategy profile set $\mathcal{X}$, whereby, at every time step $t=0,1,2,\ldots$, conditioned on the current configuration $X(t)=x$, one player $i$ is randomly selected from the player set $\mathcal{V}$ according to a distribution that assigns positive probability to all players, and she updates her action to a value $X_{i}(t+1)$ chosen uniformly from her best response set $\mathcal{B}_{i}(x_{-i})$. In fact, observe that, on the one hand, a subset of strategy profiles $\mathcal{X}^{*}\subseteq\mathcal{X}$ is BR-invariant if and only if it is a trapping set for the asynchronous best response dynamics, i.e., if $X(s)\in\mathcal{X}^{*}$ for some $s\geq 0$ implies that $X(t)\in\mathcal{X}^{*}$ for every $t\geq s$. On the other hand, $\mathcal{X}^{*}$ is globally BR-stable if and only if, for every probability distribution of the initial state $X(0)$ of the asynchronous best response dynamics, there exists a random time $T$ that is finite with probability $1$ and is such that $X(t)\in\mathcal{X}^{*}$ for every $t\geq T$.
@@
Notice that, if the action space of every player is binary, as is our case, then condition (3) implies condition (2), so that every I-path is also a BR-path, but not necessarily vice versa. As a consequence, every (globally) BR-reachable set of strategy profiles is (globally) BR-reachable. Conversely, every BR-invariant set of strategy profiles is I-invariant, but not vice versa: e.g., every non-strict Nash equilibrium is I-invariant but not BR-invariant. On the other hand, neither I-stability implies BR-stability nor vice versa: e.g., for the SNC game of Example 4 with $\alpha=0$, the set of Nash equilibria $\mathcal{N}$ is globally I-stable but not BR-stable.
@@
The SNC game on the network (a) in Figure 2 with external field $h=0$ reduces to a two-player game with binary action set $\mathcal{A}=\{\pm 1\}$ and utilities $u_{1}(x_{1},x_{2})=x_{1}x_{2}$ and $u_{2}(x_{1},x_{2})=-x_{1}x_{2}$, which is commonly referred to as the discoordination game [44].
@@
Consider a SNC game on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with two nodes $\mathcal{V}=\{1,2\}$, a single link =$\mathcal{E}=\{(1,2)\}$, and weight matrix $W=(1,0)(0,1)^{\prime}\,,$ and external field $h=(0,\alpha)$, where $\alpha$ is a scalar parameter. Then, the set of Nash equilibria is
@@
It easily verifiable that $\mathcal{N}$ is globally BR-reachable for every $\alpha$. In fact, for $\alpha\neq 0$ then the unique Nash equilibrium is strict, so that $\mathcal{N}$ is BR-invariant, hence BR-stable. On the other hand, for $\alpha=0$, there are two Nash equilibria neither of which is strict: in this case, neither $\mathcal{N}$ nor any of it subsets are BR-reachable, and there are no BR-stable sets.
@@ III-A SNC games on undirected networks
where the last identity follows from the fact that $w_{-j}=x_{-j}$ and $w_{i}=1=-x_{j}$. Similarly, we have that
@@
It then follows from [45, Corollary 2.9] that the SNC game with binary actions on $\mathcal{G}$ with external field $h$ is not an exact potential game.
(ii) If $\mathcal{G}$ is undirected, then (7) implies that $\Phi$ defined in (5) is an exact potential function for the SNC game with binary actions on $\mathcal{G}$ with external field $h$. The claim then follows from [45, Lemma 2.7].
(iii) The claim is a direct consequence of [46, Lemma 2]. ∎
@@ III-B SNC games on unsigned networks
(ii) This is a direct consequence of [40, Proposition 3(v)], which asserts that the set of Nash equilibria is globally I-stable (i.e., globally I-reachable and I-invariant) and the fact that I-reachability implies BR-reachability (c.f. Remark 2). ∎
@@
Proposition 2 ensures that every network coordination game, i.e., every SNC game with binary actions on an unsigned network $\mathcal{G}$, admits Nash equilibria and that the set $\mathcal{N}$ of Nash equilibria is globally reachable. Notice that, $\mathcal{N}$ may not be invariant, as Example 4 illustrates.
@@ III-C Signed network games on structurally balanced networks
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$, and let $\mathcal{N}$ be the set of its Nash equilibria. Given $\sigma$ in $\mathcal{X}$, consider the $[\sigma]$-transformed game and let $\mathcal{N}_{[\sigma]}$ be the set of its Nash equilibria. Then,
@@
This result coincides with [39, Lemma 1]. However, since the work [39] is developed under the “digon sign-symmetry” assumption (c.f. footnote 1), and we are working in greater generality, we present a proof in Appendix B. ∎
@@
If $\mathcal{G}$ is structurally balanced, Lemma 2 implies that there exists a gauge transformation $[\sigma]$ such that the whole node set $\mathcal{V}$ is coordinating in the $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$. It then follows from Proposition 2(ii) that the $[\sigma]$-transformed SNC game with binary action on the network $\mathcal{G}^{[\sigma]}$ with external field $h^{[\sigma]}$ has a non-empty and globally reachable set of Nash equilibria $\mathcal{N}_{[\sigma]}$. Lemma 1 then implies that the set of Nash equilibria of the SNC game with binary actions on $\mathcal{G}$ with external field $h$ satisfies $\mathcal{N}=\{[\sigma]x^{*}:\,x^{*}\in\mathcal{N}_{[\sigma]}\}\neq\emptyset$ and is globally reachable. ∎
@@
Consider the graph in Figure 4. The graph is not structurally balance. Anyway, the subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, the partition $\mathcal{R}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ with $\mathcal{R}_{1}=\{1,4\}$ and =$\mathcal{R}=\{2,3\}$ is such that $W_{\mathcal{R}_{q}\mathcal{R}_{q}}\geq 0$ for $q$ in $\{1,2\}$ and $W_{\mathcal{R}_{q}\mathcal{R}_{r}}\leq 0$ for $q\neq r$, $q$, $r$ in $\{1,2\}$. According to Proposition LABEL:pr2, there exists a gauge transformation $[\sigma]$ such that $\mathcal{R}$ is a coordinating set of $\mathcal{G}^{[\sigma]}$. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]$ is such that
@@ IV-A Existence of Nash equilibria
for some $a=\pm 1$. Our first result, stated below, guarantees the existence of such Nash equilibria in terms of two properties of the $\mathcal{R}$- and $\mathcal{S}$-restricted SNC games, respcetively: the first one, cohesiveness, limits the influence that players in $\mathcal{S}$ can have on the players in $\mathcal{R}$, the second one ensures that the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to a consensus admits a Nash equilibrium.
@@
$w_{i}^{{\mathcal{R}}}\geq 1/2w_{i}\,,\qquad\forall i\in{\mathcal{R}}\,.$
@@
Notice that the two networks of Examples 2 and 3, for which Nash equilibria did not exist, were missing just one of the two assumptions in Proposition 4. Precisely, condition (25) was not satisfied by the network in Figure 2 (a), while condition (26) was not satisfied by the network in Figure 2 (b). We also observe that global reachability is not guaranteed by conditions (25) and (26) alone, as already noticed in Example 5.
@@
$(\mathcal{N}_{[\sigma]})_{{\mathcal{S}}}^{(\tau)}=\mathcal{N}_{{\mathcal{S}}}^{(\tau)}\neq\emptyset\,.$
@@
Consider a SNC game with binary actions on $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a binary partition (22) where both $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{S}}$ are structurally balanced. If there exists $\tau$ in $\mathcal{X}_{\mathcal{R}}$ such that assumption (30) holds true, then there exists a Nash equilibrium $x^{*}\in\mathcal{N}$ satisfying equation (32).
@@
Consider a SNC game with binary actions on $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a binary partition (22) where $\mathcal{G}_{\mathcal{R}}$ is structurally balanced and $\mathcal{G}_{\mathcal{S}}$ is undirected. If there exists $\tau$ in $\mathcal{X}_{\mathcal{R}}$ such that condition (30) holds true, then there exists a Nash equilibrium $x^{*}\in\mathcal{N}$ satisfying equation (32).

### v2

@@ I Introduction
Network games comprising both coordinating and anti-coordinating interactions have been recently proposed to model the presence of anti-conformist behaviors in a social community, accounting for some form of heterogeneity [22, 23, 24, 25, 26] . More generally, games exhibiting both strategic complements and substitutes have been recently proposed in the economic literature [27] to model heterogeneous interactions, e.g., markets with coexistence of both Cournot and Bertrand type firms. Such mixed games may fail to possess Nash equilibria. A fundamental example is the matching pennies game, which is a two-player game with one coordinating and one anti-coordinating player. In contrast with the matching pennies game, one can imagine that in a scenario where most of the players are coordinating and form a ’well’ connected subset, the presence of few anti-coordinating players should not prevent the coordinating players to reach a consensus and possibly the whole system to reach a Nash equilibrium. This is one main motivation to our work.
@@ II-A Signed Network Coordination Games
Figure 1 illustrates a network with weight values reported next to the corresponding links. Notice that, for =$\mathcal{R}=\{1,2,\dots,8\}$ the subnetwork $\mathcal{G}_{\mathcal{R}}$ is unsigned, whereas for $\mathcal{S}=\mathcal{V}\setminus\mathcal{R}=\{9,10,\dots,13\}$ the subnetwork $\mathcal{G}_{\mathcal{S}}$ is undirected.
@@
We now provide an interpretation of the SNC games introduced above. The utility function (1) is the sum of two terms. The first addend in the right-hand side of (1) is a standalone term $h_{i}x_{i}$ (i.e., a term that depends only on the action of player $i$) that models the bias of player $i$ towards one of the two possible actions depending on the sign of the preference term $h_{i}$: if $h_{i}>0$ (respectively, $h_{i}<0$), then player $i$ has a bias towards action $+1$ ($-1$), in that, in the absence of the network (i.e., if $W=0$) this action would provide a utility that is $2h_{i}$ larger than its alternative $-1$ ($+1$); if $h_{i}=0$, then no bias is present in that both actions $+1$ and $-1$ would reward player $i$ with the same utility in the absence of the network. The second addend in the right-hand side of (1) is the aggregate of pairwise interaction terms $x_{i}W_{ij}x_{j}$, each of which can be interpreted as the utility of player $i$ in a two-player game with player $j$. Notice that, while the sum index $j$ in the right-hand side of (1) formally runs over all players $j$ in $\mathcal{V}$, the sum is effectively only over players $j$ such that $W_{ij}\neq 0$, i.e., out-neighbors of player $i$ in the network $\mathcal{G}$. Observe that a SNC game with binary actions is an instance of a polymatrix or pairwise-separable game [42, 43].
@@ II-B Best Responses and Nash Equilibria
The best response (BR) correspondence for a player $i$ in $\mathcal{V}$ is the set of its optimal actions given the strategy profile of the other players. Formally,
@@
The notions of BR-reachability, BR-invariance, and global BR-stability are particularly relevant for the study of the so-called asynchronous best response dynamics [44]. These are a class of asynchronous dynamics modeled as discrete-time Markov chains with finite state space coinciding with the strategy profile set $\mathcal{X}$, whereby, at every time step $t=0,1,2,\ldots$, conditioned on the current configuration $X(t)=x$, one player $i$ is randomly selected from the player set $\mathcal{V}$ according to a distribution that assigns positive probability to all players, and she updates her action to a value $X_{i}(t+1)$ chosen uniformly from her best response set $\mathcal{B}_{i}(x_{-i})$. In fact, observe that, on the one hand, a subset of strategy profiles $\mathcal{X}^{*}\subseteq\mathcal{X}$ is BR-invariant if and only if it is a trapping set for the asynchronous best response dynamics, i.e., if $X(s)\in\mathcal{X}^{*}$ for some $s\geq 0$ implies that $X(t)\in\mathcal{X}^{*}$ for every $t\geq s$. On the other hand, $\mathcal{X}^{*}$ is globally BR-stable if and only if, for every probability distribution of the initial state $X(0)$ of the asynchronous best response dynamics, there exists a random time $T$ that is finite with probability $1$ and is such that $X(t)\in\mathcal{X}^{*}$ for every $t\geq T$.
@@
Notice that, if the action space of every player is binary, as is our case, then condition (3) implies condition (2), so that every I-path is also a BR-path, but not necessarily vice versa. As a consequence, every (globally) I-reachable set of strategy profiles is (globally) BR-reachable. Conversely, every BR-invariant set of strategy profiles is I-invariant, but not vice versa: e.g., every non-strict Nash equilibrium is I-invariant but not BR-invariant. On the other hand, neither global I-stability implies global BR-stability nor vice versa: e.g., for the SNC game of Example 4 with $\alpha=0$, the set of Nash equilibria $\mathcal{N}$ is globally I-stable but not globally BR-stable.
@@
The SNC game on the network (a) in Figure 2 with external field $h=0$ reduces to a two-player game with binary action set $\mathcal{A}=\{\pm 1\}$ and utilities $u_{1}(x_{1},x_{2})=x_{1}x_{2}$ and $u_{2}(x_{1},x_{2})=-x_{1}x_{2}$. This is commonly referred to as the discoordination game [45] and known not to possess Nash equilibria.
@@
Consider a SNC game on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with two nodes $\mathcal{V}=\{1,2\}$, a single link =$\mathcal{E}=\{(1,2)\}$, and weight matrix $W=(1,0)^{\prime}(0,1)\,,$ and external field $h=(0,\alpha)$, where $\alpha$ is a scalar parameter. Then, the set of Nash equilibria is
@@
It is easily verifiable that $\mathcal{N}$ is globally BR-reachable for every $\alpha$. In fact, for $\alpha\neq 0$ then the unique Nash equilibrium is strict, so that $\mathcal{N}$ is BR-invariant, hence globally BR-stable. On the other hand, for $\alpha=0$, there are two Nash equilibria neither of which is strict: in this case, neither $\mathcal{N}$ nor any of it subsets are BR-reachable, and there are no globally BR-stable sets.
@@ III-A SNC games on undirected networks
where the last identity follows from the fact that $w_{-j}=x_{-j}$ and $w_{j}=1=-x_{j}$. Similarly, we have that
@@
It then follows from [46, Corollary 2.9] that the SNC game with binary actions on $\mathcal{G}$ with external field $h$ is not an exact potential game.
(ii) If $\mathcal{G}$ is undirected, then (7) implies that $\Phi$ defined in (5) is an exact potential function for the SNC game with binary actions on $\mathcal{G}$ with external field $h$. The claim then follows from [46, Lemma 2.7].
(iii) The claim is a direct consequence of [47, Lemma 2]. ∎
@@ III-B SNC games on unsigned networks
(ii) This is a direct consequence of [41, Proposition 3(v)], which asserts that the set of Nash equilibria is globally I-stable (i.e., globally I-reachable and I-invariant) and the fact that I-reachability implies BR-reachability (c.f. Remark 2). ∎
@@
Proposition 2 ensures that every network coordination game, i.e., every SNC game with binary actions on an unsigned network $\mathcal{G}$, admits Nash equilibria and that the set $\mathcal{N}$ of Nash equilibria is globally BR-reachable. Notice that, $\mathcal{N}$ may not be BR-invariant, as Example 4 illustrates.
@@ III-C Signed network games on structurally balanced networks
Consider a SNC game with binary actions on a network $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ with external field $h$, and let $\mathcal{N}$ be the set of its Nash equilibria. Given $\sigma$ in $\mathcal{X}$, consider the $[\sigma]$-transformed game and let $\mathcal{B}_{i}^{[\sigma]}$, for every player $i$ in $\mathcal{V}$, be its best response correspondence, and $\mathcal{N}_{[\sigma]}$ be the set of its Nash equilibria. Then,
$\mathcal{B}_{i}^{[\sigma]}(x_{-i})=\sigma_{i}\mathcal{B}_{i}(([\sigma]x)_{-i})\,,$ (17)
for every player $i$ in $\mathcal{V}$ and strategy profile $x$ in $\mathcal{X}$, and
@@
This result coincides with [40, Lemma 1]. However, since the work [40] is developed under the “digon sign-symmetry” assumption (c.f. footnote 1), and we are working in greater generality, we present a proof in Appendix B. ∎
@@
If $\mathcal{G}$ is structurally balanced, Lemma 2 implies that there exists a gauge transformation $[\sigma]$ such that the whole node set $\mathcal{V}$ is coordinating in the $[\sigma]$-transformed network $\mathcal{G}^{[\sigma]}$. It then follows from Proposition 2(ii) that the $[\sigma]$-transformed SNC game with binary action on the network $\mathcal{G}^{[\sigma]}$ with external field $h^{[\sigma]}$ has a non-empty and globally BR-reachable set of Nash equilibria $\mathcal{N}_{[\sigma]}$. Lemma 1 then implies that the set of Nash equilibria of the SNC game with binary actions on $\mathcal{G}$ with external field $h$ satisfies $\mathcal{N}=\{[\sigma]x^{*}:\,x^{*}\in\mathcal{N}_{[\sigma]}\}\neq\emptyset$ and is globally BR-reachable. ∎
@@
Consider the graph in Figure 4. The graph is not structurally balanced. Anyway, the subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, the partition $\mathcal{R}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ with $\mathcal{R}_{1}=\{1,4\}$ and =$\mathcal{R}=\{2,3\}$ is such that $W_{\mathcal{R}_{q}\mathcal{R}_{q}}\geq 0$ for $q$ in $\{1,2\}$ and $W_{\mathcal{R}_{q}\mathcal{R}_{r}}\leq 0$ for $q\neq r$, $q$, $r$ in $\{1,2\}$. According to Lemma 2, there exists a gauge transformation $[\sigma]$ such that $\mathcal{R}$ is a coordinating set of $\mathcal{G}^{[\sigma]}$. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]$ is such that
@@ IV-A Existence of Nash equilibria
for some $a=\pm 1$. Our first result, stated below, guarantees the existence of such Nash equilibria in terms of two properties of the $\mathcal{R}$- and $\mathcal{S}$-restricted SNC games, respectively: the first one, cohesiveness, limits the influence that players in $\mathcal{S}$ can have on the players in $\mathcal{R}$, the second one ensures that the $\mathcal{S}$-restricted SNC game with strategy profile of players in $\mathcal{R}$ frozen to a consensus admits a Nash equilibrium.
@@
$w_{i}^{{\mathcal{R}}}\geq w_{i}/2\,,\qquad\forall i\in{\mathcal{R}}\,.$
@@
Notice that the two networks of Examples 2 and 3, for which Nash equilibria did not exist, were missing just one of the two assumptions in Proposition 4. Precisely, condition (26) was not satisfied by the network in Figure 2 (a), while condition (27) was not satisfied by the network in Figure 2 (b). We also observe that global BR-reachability is not guaranteed by conditions (26) and (27) alone, as already noticed in Example 5.
@@
$(\mathcal{N}_{[\sigma]})_{{\mathcal{S}}}^{(\bm{1})}=\mathcal{N}_{{\mathcal{S}}}^{(\tau)}\neq\emptyset\,.$
@@
Consider a SNC game with binary actions on $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a binary partition (23) where both $\mathcal{G}_{\mathcal{R}}$ and $\mathcal{G}_{\mathcal{S}}$ are structurally balanced. If there exists $\tau$ in $\mathcal{X}_{\mathcal{R}}$ such that assumptions (30) and (31) hold true, then there exists a Nash equilibrium $x^{*}\in\mathcal{N}$ satisfying equation (33).
@@
Consider a SNC game with binary actions on $\mathcal{G}=(\mathcal{V},\mathcal{E},W)$ and a binary partition (23) where $\mathcal{G}_{\mathcal{R}}$ is structurally balanced and $\mathcal{G}_{\mathcal{S}}$ is undirected. If there exists $\tau$ in $\mathcal{X}_{\mathcal{R}}$ such that assumptions (30) and (31) hold true, then there exists a Nash equilibrium $x^{*}\in\mathcal{N}$ satisfying equation (33).
@@
Example 6 (cont’d).
Consider the graph in Figure 4. As previously observed, the subset =$\mathcal{R}=\{1,\dots,4\}$ is such that the graph $\mathcal{G}_{\mathcal{R}}$ is structurally balanced. Indeed, if we consider the gauge transformation $[\sigma]$ with $\sigma=[1,-1,-1,1,1,1]$, we obtain that $W^{[\sigma]}=[\sigma]W[\sigma]\geq 0$. Furthermore, for $h_{i}=0$, we have $w^{\mathcal{R}}|_{\mathcal{R}}=(4,3,4,3)>w^{\mathcal{S}}|_{\mathcal{R}}=(1,1,1,1)$. Since $\mathcal{G}_{\mathcal{S}}$ is undirected (no links), by Corollary 2, this implies that, for $\tau=(1,-1,-1,1)$, there exists a polarized Nash equilibrium, that is, a Nash equilibrium where $x_{i}^{*}=\tau_{i}$ for all $i$ in $\mathcal{R}$. For instance, the strategy profile $x^{*}=(1,-1,-1,1,-1,1)$ is a Nash equilibrium of the game. Notice that, since $\mathcal{G}_{\mathcal{S}}$ has no links, Corollary 1 applies too.